OpenStudy (anonymous):

given the literal ewuation xy+x/y=2+z, with y=4x+2, what is the value of z if x=1?

OpenStudy (anonymous):

I'm not sure, but i think you would take the expression that y equals (4x+2) plug that into the equation for y, THEN plug 1 in for all the x, and finally solve for z

OpenStudy (anonymous):

x(4x+2)+x/(4x+2)=2+z ----plugging in 4x+2 1(4(1)+2)+1/(4(1)+2)=2+z -----plugging in 1 for all x can you solve now?

OpenStudy (anonymous):

I should've asked, is xy+x all over y? or just x?